OpenStudy (anonymous):

Part 1: Solve the following quadratic equation using one of the methods listed below. Part 2: Using complete sentences, explain why you chose the method you used. Equation: 0 = x2 + 6x + 3 Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (anonymous):

ok, this one looks like an easy one to complete the square for, because the middle term is even and because you cant factor do you know how to start completeing the square?

OpenStudy (anonymous):

all I know if you add 9 to both sides.

OpenStudy (anonymous):

Completing the square: x^2 + 6x = -3 x^2 + 6x + 9 = -3 + 9 (x+3)^2 = 6 x + 3 = +/- sqrt6 x = -3 +/- sqrt6 \[x = -3 \pm \sqrt{6}\]

OpenStudy (anonymous):

ok, so if you add nine to both sides, you get 9 = x^2 + 6x +9 + 3 right? so you can make that 9 = (x+3)^2 +3 does that make sense?

OpenStudy (anonymous):

I chose this method because the quadratic is not factorable and is easier to use than the quadratic formula. Completing the square is a quicker way to solve quadratics like this because it doesn't require plugging into a formula like the quadratic formula.

OpenStudy (anonymous):

Thank you guys ! (:

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Could you guys help with the one I just posted?

OpenStudy (anonymous):

I'm not sure what you did after you got (x + 3)^2 = 6

OpenStudy (anonymous):

I took the square root of both sides. So then you're left with \[x + 3 = +\sqrt{6}\]

OpenStudy (anonymous):

I mean \[x + 3 = \pm \sqrt{6}\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Then you subtracted 3 to both sides?

OpenStudy (anonymous):

Yepp (:

OpenStudy (anonymous):

Okay got it (: